OpenStudy (anonymous):

if cos x = 4/11 find cos 2A (HELP PLEASE)

OpenStudy (anonymous):

by 2A do you mean 2x?

OpenStudy (anonymous):

If that's the case, then your'e going to do arc cos of 4/11, which yields x, then plug that value of x into cos2x

OpenStudy (anonymous):

i dnt get ti can you walk me through it

OpenStudy (anonymous):

ok so by 2A did you actually mean cos2x?

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

why dont you simply use this gem of a formulla : cos 2x = (2cos^2 x) - 1 hope this helps..

OpenStudy (anonymous):

no it doesnt help CAUSE I DNT KNOW HOW TO DO IT :(

OpenStudy (shubhamsrg):

nevermind,,i cant help you further then :/

OpenStudy (anonymous):

Ok so use your graphing calculator and there should be a button that's like cos, and right above it you'll see cos^-1. Plug into your calcultar, cos^-1(4/11), because what that's saying is cos of the whatever that is, is going to be 4/11. Now you have x. Now plug whatever you got for x back into your calculator.

OpenStudy (anonymous):

Back into the equation cos2x.

OpenStudy (anonymous):

wait y am i doin cos-1 instead of cos

OpenStudy (anonymous):

because that's how you're going to find x. It's just like how if you had x^2=4, you'd find the square root of both sides. In this case you're getting rid of the cos.

OpenStudy (anonymous):

so 68,67631374

OpenStudy (anonymous):

what do i do after

OpenStudy (anonymous):

ok, so is that what you're getting for cos^-1(4/11)?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok you have that value for x, now plug that in for x in cos2x. So plug in cos(2*68.67). That SHOULD be how you do it, I haven't done stuff like this in two years.

OpenStudy (anonymous):

so is it a -.7353878608

OpenStudy (anonymous):

Yeah uhhhh that should be it, apologies if it isn't though, I don't have a calculator with me.

OpenStudy (anonymous):

use to apply cos2x=cos^2x-sin^2x cox=4/11 sinx=squareroot( 1-cos^2x)